# Rethinking the effects of data contamination in Code Intelligence

### v2

@@ III-C Results of the Unpaired Contamination Setting (RQ3)
As presented in Section II-A3, unpaired contamination experiments are conducted on code translation and generation tasks for PLMs. As for LLMs, we carry out the code translation task solely for exploration. Below, we introduce and analyse the experimental results of PLMs and LLMs, respectively.
@@
Similarly, we present the PLMs’ results first, which are shown in Table IV. It should be noted that both the inputs and outputs are highlighted in bold under the Task column, as both of them are contaminated in the experimental group. As can be seen, although PLMs in the contaminated experimental groups perform slightly better than those in the uncontaminated control groups, their superiority is very limited, ranging from 0.06%–0.52% in terms of BLEU while 0.11%–0.68% in terms of METEOR, and the statistical tests still demonstrate that there is no significant difference between the two groups with p-value larger than 0.05. This suggests that merely observing both inputs and outputs independently is insufficient for the PLMs to learn their correspondence in downstream tasks, as the input-output mappings are not constructed during the pre-training stage.
@@
The experimental results of the unpaired contamination setting for LLMs are shown in Table IV, where the code translation task is conducted in both Java-to-C# and C#-to-Java translation. As can be seen, LLMs perform slightly better in the contaminated experimental group than in the uncontaminated control group across most of the translation scenarios and evaluation metrics. Nonetheless, their performance differences still cannot be considered significant, as the p-values of the statistical tests are larger than 0.05 in most cases, except the C#-to-Java translation when evaluating in the BLEU score, which has been highlighted in red. Nevertheless, their performance in terms of METEOR does not demonstrate statistically significant differences between the two groups. Besides, theoretically, unpaired contamination only involves pre-training with both input and output portions of testing sets, but the mapping between them is unlearned. Therefore, we can only infer that this is an experimental coincidence rather than contamination-induced performance overestimation.
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.531 | / / 0.676 |
NLJava | / / 0.835 | / / 0.676 | |
GPT-2 | JavaC# | / / 0.296 | / / 0.144 |
NLJava | / / 0.403 | / / 0.676 | |
LLaMA | NLJava | / / | / / |
Starcoder | NLJava | / / | / / |
III-D Results of the Paired Contamination Setting (RQ4)
@@
As illustrated in Section II-A4, paired contamination experiments are carried out on code translation and generation tasks for PLMs, while for LLMs, only the code generation task is experimented with. Below, we introduce and analyse PLMs and LLMs’ experimental results, respectively.
@@
The experimental results of PLMs are shown in Table V. As can be seen, neither the contaminated experimental group nor the uncontaminated control group manifests consistent superiority in either code translation or generation tasks and either RoBERTa or GPT-2. Besides, the statistical tests also demonstrate non-significant differences with all p-values larger than 0.05 among five consecutive trials. The performance of RoBERTa is reasonable, as its pre-training task is MLM, which is dictated to learn to infill the masked tokens during the pre-training stage and is not aligned with the inference procedure of any downstream tasks. Besides, to implement the generative task, RoBERTa need to be appended with a decoder structure, further weakening the impact of pre-training contamination on downstream tasks. On the contrary, the experimental results of GPT-2 are kind of surprising and opposite to our expectations, because it is an auto-regressive model [32] and was pre-trained by the CLM task from scratch in our experiments. Considering that the input and output portions of the testing sets are contaminated in the pre-trained corpus of GPT-2, it should have learned how to generate correct tokens during the inference stage, theoretically. But the fact is not the case. We further explore possible explanations in Section IV.
@@
As for LLMs, their experimental results are shown in Table V as well. Apparently, both LLMs’ performance in contaminated experimental groups significantly surpasses that in uncontaminated control groups. Specifically, the outperformance reaches 14.34% in terms of BLEU and 7.26% in terms of METEOR for LLaMA. As for StarCoder, the overestimation can be 7.40% and 4.07% in terms of each metric in order. Moreover, statistical tests with p-values smaller than 0.05 further testify to the significance of the difference. Considering that LLMs under test are all decoder-only architectures and are pre-trained with either CLM (i.e., LLaMA) or Fill-in-the-Middle (FIM) task (i.e., StarCoder), the explanation is straightforward. Because a decoder-only architecture implies its auto-regressive nature in practical usage. Besides, their pre-training tasks help them learn to generate subsequent tokens given prefixes, and even the FIM task broke away from the limitation of left-to-right generation [43]. Both of the above reasons make LLMs’ pre-training stage almost or exactly match their inference stage. As such, paired contamination leads to performance overestimation of LLMs in downstream code intelligence tasks.
IV Discussion
@@
As stated in RQ4, although we inject paired contaminated data into the pre-training stage of GPT-2, the overall model performance did not exhibit a significant improvement, even though it is an auto-regressive model and its pre-training task is almost the same as LLMs under test. The conclusion is completely opposite. Considering there is a fine-tuning stage between pre-training and inference for PLMs, we propose our RQ5 naturally: Whether the expected contamination effects appear if we drop the fine-tuning stage from experiments of GPT-2, and how fine-tuning influences the contamination effects?
@@
To do this, we conduct a direct inference experiment using GPT-2 with previously pre-trained models in RQ4. The experimental results are shown in Table VI. We find that the model performance in the contaminated experimental group significantly outperforms that of the uncontaminated counterparts, where the former achieves 26.52% and 20.57% of superiority in terms of BLEU and METEOR on Java-to-C# translation, respectively. As for the code generation task, contamination even brings about overestimation by 119.53% and 73.79% in terms of each metric in order. Besides, statistical tests also demonstrate the significance of the performance differences with p-values all smaller than 0.05. This indicates that contamination indeed affects GPT-2, making its performance overestimated with the pre-train-then-inference paradigm, but this effect was covered up by the subsequent fine-tuning. As mentioned in Section II-D and II-F, our fine-tuning was conducted with 10,300 and 100,000 training samples for code translation and generation tasks, respectively. Both are trained for 50 epochs. Thus, we infer that fine-tuning brings about abundant underlying knowledge and task alignment specifications [44], making the effects of contamination in the pre-training stage negligible.
@@
To further investigate the influence of fine-tuning on contamination effects, towards the code translation and generation tasks, we reduce the fine-tuning size to 500 samples and progressively enlarge the training epochs (E=0, 1, 5, 10, 25, 50) for multiple experiments, thereby controlling the learning of underlying knowledge and task alignment specifications. As with our previous RQs, we repeat the experiment 5 times for each setting, and the results are presented in Fig. 1, where epoch = 0 denotes zero-shot inference without fine-tuning. For each pair of control and experimental groups, we use a black line to connect them, thereby making their performance differences more clear. Meanwhile, p-values for each experiment are plotted with red dots connected by dashed lines in Figure 1. As can be seen, the contaminated experimental group performs significantly better than the uncontaminated counterpart initially, i.e., E10, for code translation while E5 for code generation. With the training epochs enlarging, which means more underlying knowledge and task alignment specifications are learned by GPT-2, the contamination effects gradually eliminate. Besides, we also find an interesting phenomenon: the performance of GPT-2 in both experimental and control groups shows a trend of declining first, then rising across both code tasks. A potential explanation is that the pre-training stage of our experiments mainly learn to generate Java code given Java prefixes, which is relatively different from the fine-tuning tasks, i.e., code translation and generation. Thus, there are certain conflicts between them, leading to the re-learning of the task alignment on both downstream code tasks at the initial fine-tuning stage. With the training continuing, GPT-2 gradually overcomes the task alignment conflicts, thereby recovering and even improving its performance.
Task | BLEU (Experimental / Control / p-value) | METEOR (Experimental / Control / p-value) |
JavaC# | / / | / / |
NLJava | / / | / / |
V Implications
This paper reveals the impact of various data contamination settings on the performance of PLMs and LLMs in practical usage. Based on the experimental results, we summarise the following implications.
V-A Implications for Researchers
This paper serves as the first systematic study delving into the practical effects of different contamination settings on PLMs and LLMs. To achieve this, we design rigorous experiments for both PLMs and LLMs. In particular, we propose a series of efficient and reproducible methods to extract contaminated samples from tremendous pre-training corpora of LLMs and construct uncontaminated counterparts applicable for diverse contamination settings, which can be reused or refined by future related research. Secondly, our research concluded that contamination effects do not appear in all scenarios and all models. Researchers can conduct their LLM/PLM-related experiments according to our conclusions to deal with the data contamination problem more precisely.
V-B Implications for Practitioners
@@
Based on our extensive experimental results, we provide guidance for developers to better understand the impact of data contamination in code intelligence. Firstly, we find that although PLMs suffer from contamination during pre-training, their performance does not show significant improvement on the test set after fine-tuning, indicating that the impact of contamination can be effectively mitigated to some extent. Therefore, deploying PLMs with fine-tuning in practice is safe for data contamination, contrary to the conclusion from NLP research [45]. Therefore, contamination in fine-tuning should be given more attention instead of pre-training for PLMs. Secondly, practitioners must exercise heightened vigilance around paired contamination for the usage of LLMs, as this configuration makes model performance significantly overestimated. This differential impact between PLMs and LLMs necessitates tailored contamination mitigation strategies in practice.
VI Threats to Validity
@@
Limited contamination scale: In this study, the proportion of contaminated data in experiments of both PLMs and LLMs is relatively small compared to their pre-training corpora, without further exploration of the larger scale of contamination. However, it is worth noting that in real-world scenarios, benchmark evaluation datasets, such as HumanEval [15] and ClassEval [46], typically constitute only a tiny fraction compared to the scale of the PLM/LLMs’ pre-training corpora, which aligns with our experimental settings. Therefore, our experimental results remain reasonably credible for code intelligence practice despite this potential limitation.
@@
Perturbation quality: when constructing uncontaminated datasets for LLMs, we manually perturb the contaminated data to be unseen to LLMs, leading to two threats concerning data quality, i.e., 1) whether the perturbed sample can be considered uncontaminated, 2) whether the perturbation affects the sample complexities. To achieve this, we adopt a relatively high perturbation rate of 14.19%–48.58% on average across various experiments (as shown in Table I), thereby making uncontaminated samples as different from the original versions as possible, thereby alleviating the former threat. Besides, two authors of this paper participate in the perturbation with a double-checking mechanism to ensure the consistency of token lengths and cyclomatic complexity before and after perturbation. Therefore, the threat of the latter is also minimised.
VII Related Work
VII-A Code Intelligence Tasks on PLMs and LLMs

### v3

@@ 3.3. Results of the Unpaired Contamination Setting (RQ3)
As presented in Section 2.1.3, unpaired contamination experiments are conducted on all code tasks for PLMs, while LLMs are only evaluated on the code translation task solely. Below, we introduce and analyze the experimental results of PLMs and LLMs, respectively.
@@
Similarly, we first present the results of PLMs (Pre-train, fine-tune, then infer), which are shown in Table 4. As can be seen, although PLMs in the contaminated experimental groups perform slightly better than those in the uncontaminated control groups on the code translation and generation tasks overall, their superiority is very limited, ranging from (-0.15%)–0.42% in terms of BLEU while 0.01%–0.40% in terms of METEOR, across both PLMs. As for the code summarization task, the performance comparison is totally opposite, where the PLMs mostly perform weaker in the contaminated experimental groups. But, the performance discrepancies are still neglectable, on average lagging 0.13%–0.03% and 0.30%–0.33% in terms of each metric across both PLMs. Moreover, the statistical tests also demonstrate that there is no significant difference between the two groups above with p-value larger than 0.05. This suggests that merely observing both inputs and outputs independently is insufficient for the PLMs to learn their correspondence in downstream tasks, as the input-output mappings are not constructed during the pre-training stage.
@@
The experimental results of the unpaired contamination setting for LLMs (Pre-train, then infer) are shown in Table 4. As can be seen, neither the contaminated experimental group nor the uncontaminated control group demonstrates significant superiority over the other one on the code translation task. The performance variation fluctuates between 0.59%–1.56% in terms of BLEU and (-0.92%)–0.64% in terms of METEOR across both LLMs on average. Meanwhile, the above performance differences still cannot be considered significant, as the p-values of the statistical tests are larger than 0.05. The above experimental results are rational because, theoretically, unpaired contamination only involves pre-training with both input and output portions of testing sets, but the mapping between them is unlearned.
| Metric | Model Performance (Experimental Group / Control Group / p-value) | |||
| RoBERTa | GPT-2 | LLaMA | StarCoder | |
| Code Translation: Java C# | ||||
| BLEU | 86.18±0.20 / 86.11±0.09 / 0.531 | 37.43±0.09 / 37.16±0.36 / 0.296 | —— / —— / —— | —— / —— / —— |
| METEOR | 88.93±0.16 / 88.99±0.16 / 0.676 | 48.01±0.03 / 47.75±0.32 / 0.144 | —— / —— / —— | —— / —— / —— |
| Code Translation: Python Java | ||||
| BLEU | 70.09±0.51 / 69.78±0.22 / 0.346 | 48.74±0.10 / 48.67±0.13 / 0.402 | —— / —— / —— | —— / —— / —— |
| METEOR | 70.39±0.54 / 69.97±0.22 / 0.249 | 49.61±0.20 / 49.57±0.12 / 0.531 | —— / —— / —— | —— / —— / —— |
| Code Generation: NL Java | ||||
| BLEU | 40.86±0.50 / 40.77±0.43 / 0.835 | 17.55±0.32 / 17.72±0.20 / 0.403 | 35.16±1.56 / 30.75±1.24 / 0.012 | 35.27±0.73 / 32.84±0.67 / 0.012 |
| METEOR | 51.33±0.48 / 51.47±0.38 / 0.676 | 21.95±0.48 / 21.98±0.13 / 0.676 | 49.67±1.50 / 46.31±1.37 / 0.012 | 50.86±0.54 / 48.87±0.74 / 0.012 |
| Code Generation: NL Python | ||||
| BLEU | 37.65±0.23 / 37.39±0.13 / 0.059 | 15.20±0.27 / 15.47±0.09 / 0.095 | 7.21±0.26 / 6.59±0.28 / 0.012 | 9.27±0.51 / 7.93±0.24 / 0.012 |
| METEOR | 45.60±0.26 / 45.31±0.09 / 0.095 | 31.17±0.41 / 31.63±0.15 / 0.144 | 18.57±0.31 / 17.53±0.19 / 0.012 | 21.27±0.37 / 19.95±0.46 / 0.012 |
| Code Summarization: Java NL | ||||
| BLEU | 39.06±0.60 / 38.68±0.42 / 0.676 | / 2.80±0.06 / 0.834 | 10.11±0.11 / 8.77±0.28 / 0.012 | 11.60±0.30 / / 0.012 |
| METEOR | 48.65±0.61 / 48.51±0.34 / 0.295 | / 18.72±0.22 / 0.835 | 39.44±0.17 / 37.45±0.38 / 0.012 | 43.50±0.54 / / 0.012 |
| Code Summarization: Python NL | ||||
| BLEU | / 37.09±0.25 / 0.095 | 7.17±0.03 / 7.15±0.03 / 0.598 | 7.50±0.28 / 6.55±0.15 / 0.012 | / 6.41±0.13 / 0.012 |
| METEOR | 46.73±0.40 / 46.32±0.16 / 0.095 | / 22.12±0.10 / 0.142 | 34.11±0.35 / 31.02±0.19 / 0.012 | 22.54±0.24 / 21.59±0.31 / 0.012 |
3.4. Results of the Paired Contamination Setting (RQ4)
@@
As illustrated in Section 2.1.4, paired contamination experiments are carried out on all three code tasks for PLMs, while for LLMs, only the code generation and summarization tasks are experimented with. Below, we introduce and analyze PLMs’ and LLMs’ experimental results, respectively.
@@
Regarding the PLMs (Pre-train, fine-tune, then infer), the experimental results are shown in Table 5. Overall, for the code translation task, experimental groups with data contamination slightly outperform the control group for both RoBERTa and GPT-2, with average improvements of 0.26%–0.44% in BLEU and 0.27%–0.31% in METEOR. In contrast, on the code generation and summarization tasks, the two PLMs exhibit divergent performance trends. For RoBERTa, the experimental groups perform slightly better against the corresponding control groups by 0.76% and 0.39% in terms of BLEU and METEOR on average, respectively. Nevertheless, the overall results are opposite for GPT-2, where the experimental groups slightly lag on average by 0.61% and 0.51% in terms of each metric in order. Even though they have diverse performance discrepancies between the two groups, the statistical tests still reflect non-significant differences for all the above results, with p-values all larger than 0.05. The limited contamination effect towards RoBERTa is reasonable, as its pre-training task is MLM, which is designed to learn to infill the masked tokens during the pre-training stage and is not aligned with the inference procedure of any downstream tasks. Besides, to implement the generative task, RoBERTa needs to be appended with a decoder structure, further weakening the impact of pre-training contamination on downstream tasks. On the contrary, the experimental results of GPT-2 are kind of surprising and opposite to our expectations, because it is an auto-regressive model (Radford et al., 2018) and was pre-trained by the CLM task from scratch in our experiments. Considering that the input and output portions of the testing sets are contaminated in the pre-trained corpus of GPT-2, it should have learned how to generate correct tokens during the inference stage, theoretically. But the fact is not the case. We further explore possible explanations in Section 3.5.
@@
As for LLMs (Pre-train, then infer), their experimental results are shown in Table 5 as well. Apparently, both LLMs’ performance in contaminated experimental groups significantly surpasses that in uncontaminated control groups. Specifically, the outperformance reaches 9.41%–15.28% in terms of BLEU and 5.31%–9.96% in terms of METEOR for LLaMA. As for StarCoder, the overestimation can be 7.40%–16.94% and 4.07%–15.57% in terms of each metric in order. Moreover, statistical tests with p-values smaller than 0.05 further testify to the significance of the difference. Considering that LLMs under test are all decoder-only architectures and are pre-trained with either CLM (i.e., LLaMA) or Fill-in-the-Middle (FIM) task (i.e., StarCoder), the explanation is straightforward. Because a decoder-only architecture implies its auto-regressive nature in practical usage. Besides, their pre-training tasks help them learn to generate subsequent tokens given prefixes, and even the FIM task broke away from the limitation of left-to-right generation (Bavarian et al., 2022). Both of the above reasons make LLMs’ pre-training stage almost or exactly match their inference stage, thereby leading to performance overestimation of LLMs in downstream code intelligence tasks.
3.5. Discussion Experiments
@@
As stated in RQ4, although we inject paired contaminated data into the pre-training stage of GPT-2, the overall model performance did not exhibit a significant improvement, even though it is an auto-regressive model and its pre-training task is almost the same as LLMs under test. The conclusion is completely opposite. Considering there is a fine-tuning stage between pre-training and inference for PLMs, we naturally propose our RQ5: Whether the expected contamination effects appear if we drop the fine-tuning stage from experiments of GPT-2, and how fine-tuning influences the contamination effects?
@@
To do this, we conduct a direct inference experiment using GPT-2 with previously pre-trained models in RQ4. The experimental results are shown in Table 6. We find that the model performance in the contaminated experimental group significantly outperforms that of the uncontaminated counterparts by 26.52%–47.00% and 20.57%–34.73% in terms of BLEU and METEOR on the code translation task, respectively. As for the code generation task, contamination even brings about overestimation by 77.48%–119.53% and 46.98%–73.79% in terms of each metric in order. Besides, GPT-2 also suffers an overestimation on the code summarization task by 17.86%–36.14% and 22.77%–35.50% in each metric. Besides, statistical tests also demonstrate the significance of the performance differences with p-values all smaller than 0.05. This indicates that contamination indeed affects decoder-only PLMs (e.g., GPT-2), regardless of static- (e.g., Java) or dynamic-typed PLs (e.g., Python), making its performance overestimated with the pre-train-then-inference paradigm, but this effect was covered up by the subsequent fine-tuning. As mentioned in Section 2.4 and 2.6, our fine-tuning was conducted with 100,000 training samples at most for each code task, and was trained for 50 epochs, respectively. Thus, we infer that fine-tuning brings about abundant underlying knowledge and task alignment specifications (Zheng et al., 2025), making the effects of contamination in the pre-training stage negligible.
@@
To further investigate the influence of fine-tuning on contamination effects, towards each studied code task, we reduce the fine-tuning size to 500 samples and progressively enlarge the training epochs (E=0, 1, 5, 10, 25, 50) for multiple experiments, thereby simulating the expansion of the fine-tuning scale and controlling the extent of learning underlying knowledge and task alignment specifications. As with our previous RQs, we repeat the experiment 5 times for each setting, and the results are presented in Fig. 1, where E = 0 denotes zero-shot inference without fine-tuning. For each pair of control and experimental groups, we use a black line to connect them, thereby making their performance differences clearer. Meanwhile, p-values for each experiment are plotted with red dots connected by dashed lines in Figure 1. As can be seen, the contaminated experimental group overall performs significantly better than the uncontaminated counterpart initially (e.g., E5 for almost all settings). With the training epochs enlarging, which means more underlying knowledge and task alignment specifications are learned by GPT-2, the contamination effects gradually eliminate. Besides, we also find an interesting phenomenon: the average performance in BLEU and METEOR of GPT-2 under both experimental and control groups shows a trend of declining first, then rising across almost all code tasks. A potential explanation is that the pre-training stage of our experiments mainly learn to generate Java/Python code given Java/Python prefixes, which is relatively different from the fine-tuning tasks, i.e., code translation, generation, and summarization. Thus, there are certain conflicts between them, leading to the re-learning of the task alignment on all downstream code tasks at the initial fine-tuning stage. With the training continuing, GPT-2 gradually overcomes the task alignment conflicts, thereby recovering and even improving its performance, which is also a phenomenon revealed in (Zheng et al., 2025).
| Task | In / Out | Model Performance (Experimental Group / Control Group / p-value) | |
| BLEU | METEOR | ||
| Code Translation | JavaC# | 22.47±0.74 / 17.76±0.15 / 0.012 | 35.76±0.80 / 29.66±0.23 / 0.012 |
| PythonJava | 29.43±0.08 / 20.02±0.11 / 0.012 | 36.89±0.26 / 27.38±0.22 / 0.012 | |
| Code Generation | NLJava | 4.72±0.86 / 2.15±0.04 / 0.012 | 9.35±1.41 / 5.38±0.52 / 0.012 |
| NLPython | 7.17±0.23 / 4.04±0.22 / 0.012 | 21.93±0.23 / 14.92±0.31 / 0.012 | |
| Code Summarization | JavaNL | 1.13±0.02 / 0.83±0.03 / 0.012 | 13.32±0.08 / 9.83±0.13 / 0.012 |
| PythonNL | 2.64±0.17 / 2.24±0.11 / 0.012 | 14.99±0.28 / 12.21±0.28 / 0.012 | |
4. Implications
Implications for Researchers: This paper serves as the first systematic study delving into the practical effects of different fine-grained contamination settings on diverse PLM/LLMs, PLs, and tasks in the domain of code intelligence. When exploring data contamination for LLMs, the experimental/control group datasets construction method we propose in Section 2.5 is systematic against the massive LLMs’ pre-training corpora (¿ 1 trillion tokens), and is reproducible for future related studies. Secondly, considering that code corpora of most LLMs are extracted from Code Hosting Platforms (CHPs), e.g., GitHub, for tasks composed of obvious unpaired/partially-occurred samples from CHPs, e.g., test case generation and code translation tasks, researchers need not continuously build new benchmarks based on model training cut-off date to avoid the bias brought by data contamination. Our work provides empirical evidence to support this and reduces the burden of such labor-intensive efforts. Thirdly, for data contamination mitigation studies, our work reveals that researchers should focus on the fine-tuning phase if the PLM was fine-tuned before inference. However, for the pretrain–inference paradigm, the usage scenario of most LLMs, the pretraining stage should be concentrated.
@@ 4. Implications
Implications for Practitioners: Firstly, in Section 3.5, we find that although PLMs suffer from contamination during pre-training, their performance does not show significant improvement on the test set after fine-tuning, indicating that the impact of contamination can be effectively mitigated to some extent. Therefore, deploying PLMs with fine-tuning in practice is safe for data contamination of the pre-training stage, contrary to the conclusion from NLP research (Jiang et al., 2024b) mentioned in Section 2.1.1. In other words, contamination in fine-tuning should be given more attention instead of pre-training for PLMs. Secondly, when LLM practitioners are doing data cleaning before LLM pre-training (Touvron et al., 2023; Jiang et al., 2025), they need not allocate efforts to code samples that are unpaired or only partially occur in follow-up benchmarks, because these kinds of contamination carry insignificant effects on downstream performance, as we revealed in Section 3.
5. Threats to Validity
@@
Limited contamination scale: In this study, the proportion of contaminated data in experiments of both PLMs and LLMs is relatively small compared to their pre-training corpora, without further exploration of the larger scale of contamination. However, it is worth noting that in real-world scenarios, benchmark evaluation datasets, such as HumanEval (Chen et al., 2021) and ClassEval-T (Xue et al., 2025), contain only 164 Python functions and 282 class-level code pairs, aligning with our experimental settings on data volumes. Therefore, our experimental results remain reasonably credible for code intelligence practice despite this potential limitation.
@@
Perturbation quality: When constructing uncontaminated datasets for LLMs, we manually perturb the contaminated data to be unseen to LLMs, leading to two threats concerning data quality: 1) whether the perturbed sample can be considered uncontaminated, and 2) whether the perturbation affects the sample complexities. Following most previous workarounds for avoiding data contamination (Rao et al., 2025; Wang et al., 2023a; Sallou et al., 2024), we adopt a relatively high perturbation rate of 20.38%–24.65% on average across various experiments, shown in Table 1, to craft LLM-unseen samples. Considering the massive scale of LLMs’ pretraining corpora (¿ 1 trillion tokens), we admit the above procedure cannot absolutely ensure that all perturbed code is uncontaminated, but it makes approximately uncontaminated testing sets. Compared with the ideal uncontaminated testing sets that nearly cannot be reached, our constructed uncontaminated testing sets can obtain the same conclusion, just with differences in the degrees of significance. As for the second threat, two authors of this paper participate in the perturbation with a double-checking mechanism to ensure the consistency of token lengths and cyclomatic complexity before and after perturbation. According to the dataset statistics listed in Table 1, this threat is also minimized.
6. Related Work
6.1. Code Intelligence Tasks on PLMs and LLMs